40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 31 23:12:17 2018, gautam, Update, LSC, DRMI locking attempt REFL55_whtCheck.pdf
    Reply  Fri Jun 1 01:22:50 2018, gautam, Update, LSC, DRMI locking restored DRMI_20180531.png
       Reply  Fri Jun 8 01:58:56 2018, gautam, Update, LSC, DRMI locking attempt again 
          Reply  Tue Jun 12 03:22:25 2018, gautam, Update, LSC, AUX laser shuttered 
             Reply  Wed Jun 13 01:02:40 2018, gautam, Update, LSC, Reliable and repeatable 1f DRMI locking 
                Reply  Wed Jun 13 11:17:40 2018, gautam, Update, LSC, PRCL loop shape anomaly PRCL_12Jun2018_WeirdShape.pdf
                   Reply  Thu Jun 14 00:40:42 2018, gautam, Update, LSC, PRCL loop shape anomaly 
                Reply  Thu Jun 14 18:09:24 2018, gautam, Update, LSC, Reliable and repeatable 1f DRMI locking DRMI1f_June14.pdf
                   Reply  Fri Jun 15 00:53:21 2018, gautam, Update, LSC, Calibrated MICH spectrum C1NB_disp_40m_MICH_NB_2018-06-14.pdf
Message ID: 13933     Entry time: Fri Jun 8 01:58:56 2018     In reply to: 13908     Reply to this: 13948
Author: gautam 
Type: Update 
Category: LSC 
Subject: DRMI locking attempt again 

Given the various changes to the IFO config since last Thursday when I was last able to lock the DRMI, I wanted to try once again tonight. However, I had no success. By my judgement, the alignment is fine as judged by looking at mode flashes on the cameras. However, despite following the usual alignment procedures, I did not get a single lock in tonight. indecision

Perhaps we can use a flip mount on the BS that combines the PSL and AUX beams on the AS table, so we have the option of recovering the usual IFO config when we so desire - while Jon needs the SRC locked for his measurement, it would be nice to not have to figure out the correct demod phases etc each time there is a change in the optical setup of the AUX beam.

ELOG V3.1.3-